-4x -y+2 (x-3y) ; x=3 , y=-2
Answer by oberobic(2304)

-4x -y+2 (x-3y) ; x=3 , y=-2
.
substitute the values
.
-4(3) -(-2) +2(3) -6(-2) = -12 +2 +6 -12 = -16
.
Done.